1 Reply Latest reply on Feb 22, 2019 3:08 PM by mnevill

    BLACK NPD BEAD signal output unstabil

    cocones

      Hi all,

      I all way use white bead for NPD dedectors. When I change to Black bead. ( applyed replacement procedure)

      Signal output unstable.

      set to 30 pA. start adjust, 30min later complet adjust . But output goes to 15 pA and up to 30 pA. this cycle all ways continuous

      h2 : 3 ml 

      air: 60 ml

      h2 supplay pressure : 1.5 bar

      air supply pressure: 3  bar

      bead voltage off position : output 2.1pA.

      I waiting your advice.

       

      Best regards.

       

       

      @